OpenStudy (anonymous):

1 + 1/3 = (3/3) + (1/3)

OpenStudy (anonymous):

yes, yes it does

OpenStudy (anonymous):

THE ANSWER????

OpenStudy (anonymous):

there is no answer. You're not solving for anything. It's just a statement

OpenStudy (anonymous):

Also one more Maria and paul collect marbles. Paul has one third as many marbles as maria. Together they have 48 marbles . How many marbles does Maria have?

OpenStudy (anonymous):

p+m=48 p=1/3m 1+1/3m=4/3m 48*3/4=36

OpenStudy (anonymous):

How many does Maria Have?

OpenStudy (anonymous):

maria has 36. I guess i didnt make it to clear. p=number paul has and m= number maria has. so 4/3m=48 m=48/(4/3) or m=48*(3/4) m=36